10 years before,dad's age was 3 times his son's age,10 years after,dad's age is double of son's age.what's son's age now?
Answer by ankor@dixie-net.com(22740)

10 years before,dad's age was 3 times his son's age,
10 years after,dad's age is double of son's age.
what's son's age now?
:
let d = dad's present age
let s = son's present age
:
Write an equation for each statement.
:
"10 years before, dad's age was 3 times his son's age,"
d - 10 = 3(s-10)
d - 10 = 3s - 30
d = 3s - 30 + 10
d = 3s - 20
:
"10 years after, dad's age is double of son's age."
d + 10 = 2(s+10)
d + 10 = 2s + 20
d = 2s + 20 - 10
d = 2s + 10
:
what's son's age now?
in the above equation, replace d with (3s-20)
3s - 20 = 2s + 10
3s - 2s = 10 + 20
s = 30 yrs old is the son's present age
:
:
You can find dad's age and check solutions in both statements